OpenStudy (anonymous):

Find the exact area of the region that lies under the curve given by y=2sin(x)−sin(2x),0≤x≤π

OpenStudy (anonymous):

I know to change the problem into an integral notation \[\int\limits_{0}^{\pi} 2\sin(x)-\sin(2x)\] u=2x du=2 dx dx=1/2 du the question is what do i do with the other x?

zepdrix (zepdrix):

If you want to, you can split up the integral, and only apply the `u sub` to the second one.\[\int\limits\limits_{0}^{\pi} 2\sin(x)-\sin(2x) \quad = \quad \int\limits\limits_{0}^{\pi} 2\sin(x)dx-\int\limits\limits_{0}^{\pi} \sin(2x)dx\]

OpenStudy (anonymous):

why not the first one too?

zepdrix (zepdrix):

It doesn't look like you need to apply it to the first one :O unless it's confusing you.

OpenStudy (anonymous):

and make the u=x and du=1 dx and dx= du

OpenStudy (anonymous):

yes it is confusing me a lot

zepdrix (zepdrix):

Sure you can do that if you want :) Two separate u-subs so both answers are in terms of u. That will work just fine ^^

OpenStudy (anonymous):

ok ill try that thanks for your help

zepdrix (zepdrix):

When you get more comfortable with integrals, you'll be able to skip the `u sub` on problems like this. Because it's really just thinking of the chain rule in a different way. If we wanted to take the `derivative` of \(\sin(2x)\) a \(2\) would pop out due to the chain rule right? Well, since we're integrating it, the opposite will happen, a \(\dfrac{1}{2}\) will pop out, so to speak :D

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi} 2 \sin(u)du-\int\limits_{0}^{\pi}\sin(u)\frac{ 1 }{2 }du\]

zepdrix (zepdrix):

yep looks good.

OpenStudy (anonymous):

now should i combine them so they are 2sin(u)-sin(u)1/2 du?????

OpenStudy (anonymous):

and the 2's cancel

zepdrix (zepdrix):

Well we can't combine them since it's a `definite integral` and we applied two DIFFERENT substitutions. They now have different limits of integrations.\[\huge \int\limits\limits_{x=0}^{\pi} 2 \sin(u)du-\int\limits\limits_{x=0}^{\pi}\sin(u)\frac{ 1 }{2 }du\]

zepdrix (zepdrix):

See how our limits are in terms of `x`? We need limits in terms of `u`.

OpenStudy (anonymous):

i got there

OpenStudy (anonymous):

I just don't know what to do next

zepdrix (zepdrix):

I think the limits would change on the second one yes?\[\huge \int\limits\limits\limits_{x=0}^{\pi} 2 \sin(u)du-\int\limits\limits\limits_{\color{royalblue}{u=0}}^{\color{royalblue}{2\pi}}\sin(u)\frac{ 1 }{2 }du\]

zepdrix (zepdrix):

This is a really bad way to do this problem :( Hmmm.

zepdrix (zepdrix):

We really should start over and do it differently. But if you want to continue doing it this way, then that's ok, we can write the integrals like this, pulling the constants outside.\[\huge 2\int\limits\limits_{x=0}^{\pi} \sin(u)du-\frac{1}{2}\int\limits\limits_{\color{royalblue}{u=0}}^{\color{royalblue}{2\pi}}\sin(u)du\]Then we can integrate them separately from here.

zepdrix (zepdrix):

Woops, that first one should be u=0 at the bottom. my bad.

zepdrix (zepdrix):

Uh oh did I confuse you? Is your brain gonna esplode? :c

zepdrix (zepdrix):

I hate when that happens... brains everywhere...

OpenStudy (anonymous):

no not yet im trying to finish the problem

OpenStudy (anonymous):

haha nope not me

OpenStudy (anonymous):

is the right answer -3?

zepdrix (zepdrix):

I think the answer is +4. Lemme check my work again.

zepdrix (zepdrix):

Yah it's 4. Did we run into a little mistake somewhere? <:o

OpenStudy (anonymous):

yes i have an idea to where that happened thanks for the help

zepdrix (zepdrix):

oh ok cool c: